# Ultra-High-Energy Cosmic Rays from Neutrino-Emitting Tidal Disruption Events

### v2

@@ Abstract
We revisit the Ultra-High-Energy Cosmic Ray (UHECRs) production in Tidal Disruption Events (TDEs) in the light of recent neutrino-TDE associations. We use an isotropically emitting source-propagation model, which has been developed to describe the neutrino production in AT2019dsg, AT2019fdr, and AT2019aalc. These TDEs have strong dust echoes in the infrared range, which are potentially linked with the neutrino production. A mechanism where neutrinos originate from cosmic ray scattering on infrared photons implies cosmic rays in the ultra-high energy range, thus suggesting a natural connection with the observed UHECR. We extrapolate the three TDE associations to a population of neutrino- and UHECR-emitting TDEs, and postulate that these TDEs power the UHECRs. We then infer the source composition, population parameters, and local rates that are needed to describe UHECR data. We find that UHECR data point towards a mix of light to mid-heavy injection isotopes, which could be found, e.g., in oxygen-neon-magnesium white dwarfs, and to a contribution of at least two groups of TDEs with different characteristics, dominated by AT2019aalc-type events. The required local TDE rates of , however, are more indicative of the disruption of main sequence stars. We propose an enhanced efficiency in the acceleration of heavier nuclei that could address this discrepancy. The predicted diffuse neutrino fluxes suggest a population of astrophysical neutrino sources that can be observed by future radio neutrino detection experiments.
@@ 1 Introduction
Tidal Disruption Events (TDEs) are astrophysical transients originating from stars that are tidally destroyed by a supermassive black hole (SMBH). The accretion of bound stellar materials could power a month to year-long flare (Rees, 1988; Phinney, 1989), whose spectrum extends from radio to X-ray frequencies. So far, TDEs and TDE candidates have been identified by the Zwicky Transient Facility (ZTF, Bellm et al., 2019), the Wide-field Infrared Survey Explorer (WISE, Wright et al., 2010), and X-ray/radio surveys, such as eROSITA (Sazonov et al., 2021) and Very Large Array Sky Survey (VLASS, Lacy et al., 2020). Remarkably, the correlation study of an 200 TeV IceCube neutrino event IC-191001A and optical transients has revealed that a TDE, AT2019dsg, could be its most likely astrophysical counterpart (Stein et al., 2021). The subsequent joint analyses of IceCube neutrino events and electromagnetic signals have discovered more TDE candidates with potential neutrino associations, including AT2019fdr (Reusch et al., 2022), AT2019aalc (van Velzen et al., 2021), two dust-obscured TDE candidates (Jiang et al., 2023), and AT2021lwx (located at a relatively high redshift , Yuan et al., 2024a).
@@
The potential correlations between TDEs and astrophysical neutrinos provide the smoking-gun evidence of cosmic ray accelerations since high-energy neutrinos are generated from the interactions between energetic nuclei and ambient radiation/matter backgrounds. However, for TDEs, the sites for particle accelerations and interactions are still unclear. Many models including relativistic jets (Wang et al., 2011; Wang & Liu, 2016; Dai & Fang, 2017; Lunardini & Winter, 2017; Senno et al., 2017), accretion disks (Hayasaki & Yamazaki, 2019), wide-angle outflows/hidden winds (Fang et al., 2020), and tidal stream interactions (Dai et al., 2015; Hayasaki & Yamazaki, 2019) have been proposed as the origin of the non-thermal electromagnetic (EM) and neutrino emission from TDEs, which could potentially explain these TDE-neutrino coincidences (e.g., Liu et al., 2020; Murase et al., 2020; Winter & Lunardini, 2021; Wu et al., 2022; Hayasaki, 2021; Winter & Lunardini, 2023; Yuan & Winter, 2023).
@@
The non-detection of -rays from these neutrino-coincident TDEs by the Fermi Large Space Telescope (LAT) exerts stringent constraints on the TDE source parameters. In particular, the recent investigation of electromagnetic radiation induced by secondary particles of hadronic (e.g., and ) processes demonstrates that the production of Ultra-High-Energy Cosmic Rays (UHECRs) with energies eV is permitted in an isotropic radiation region of radius cm without exceeding the -ray upper limits (Yuan & Winter, 2023). Moreover, it was observed that the neutrino-TDE associations are accompanied by strong dust echoes in the infrared wavelength band, and that the arrival times of the coincident neutrinos were close to the peak of the infrared luminosity (which is delayed with respect to the OUV blackbody peak). This led to proposing cosmic ray scattering on infrared photons as a primary neutrino production mechanism, where cosmic-ray primary energies in the UHECR range are needed to exceed the photopion production threshold (e.g., Winter & Lunardini, 2023; Yuan & Winter, 2023; Yuan et al., 2024a). These conclusions support the idea that TDEs can be promising UHECR sources, and thus offer a possible answer to the long-standing problem of the origin of the observed UHECR flux (Farrar & Gruzinov, 2009; Farrar & Piran, 2014; Zhang et al., 2017; Guépin et al., 2018; Biehl et al., 2018b; Piran & Beniamini, 2023).
@@
The origin of UHECRs is another puzzle in astroparticle physics (Alves Batista et al., 2019). It is challenging to associate the measured CRs with a specific source class, (Capel & Mortlock, 2019; Abdul Halim et al., 2024a; Farrar, 2014; Shaw et al., 2022). Even with theoretical modeling of promising source classes, it remains unclear which specific source class is responsible for UHECRs, including Gamma-Ray Bursts (GRBs, Heinze et al., 2020; Boncioli et al., 2019; Liu et al., 2011; Murase et al., 2006; Samuelsson et al., 2019; Vietri, 1995; Waxman, 1995) , Active Galactic Nuclei (AGNs, Rodrigues et al., 2018; Resconi et al., 2017; Dermer & Razzaque, 2010), Starburst galaxies (Romero, G. E. et al., 2018; Anchordoqui, 2018; Anchordoqui et al., 1999), and Galaxy clusters (Kang et al., 1997; Murase et al., 2008; Kotera et al., 2009). Sources that produce UHECRs also might emit neutrinos, which contribute to the high-energy diffuse neutrino background. Therefore, the joint consideration of UHECR data and the IceCube diffuse neutrino background would provide valuable insights into the physical environments (e.g., the abundance of intermediate and heavy elements), the acceleration mechanisms (e.g., how the acceleration efficiency depends on the nuclei mass), and redshift distributions of UHECR sources.
@@
In this study, we revisit the potential role of TDEs as sources of UHECRs, motivated by the recent identification of a possible link between three observed TDEs (AT2019dsg, AT2019fdr, and AT2019aalc) and high-energy neutrino detections by IceCube. Based on the models presented in Winter & Lunardini (2023), we aim to explore the range of parameters that describe TDEs by analyzing various factors, including the type of disrupted stars, the size of the production region, and the maximum energies of CRs. By examining these parameters, we aim to identify scenarios that reproduce the UHECR data and satisfy bounds from the diffuse neutrino flux observations. To achieve this, we develop detailed computational models of cosmic ray emission in TDEs using the NeuCosmA (Baerwald et al., 2011; Hümmer et al., 2010) software and then model the transport of UHECRs from sources to Earth using the PriNCe code (Heinze et al., 2019). By comparing the results with the Akaike Information Criterion, we determine which scenarios best fit current observational data, leading to an improved understanding of TDEs as potential sources of UHECRs.
@@
The paper is structured as follows: In Sec. 2, we develop a model of TDEs using AT2019dsg, AT2019fdr, and AT2019aalc as templates. We identify a set of parameters that will be used in a fit to the data, and determine the ranges of these parameters that are allowed by the observations of AT2019dsg, AT2019fdr, and AT2019aalc. In Sec. 3, we review the current understanding of TDEs as a population of sources, and introduce our scenarios of reference of the UHECR-neutrino-emitting population of TDEs. Using a fitting procedure, we compare these scenarios with the data on the UHECR spectrum and composition from Auger, and present a detailed analysis of the best-fit scenario and how well it compares to the data. In Sec. 5 we elaborate on the question of the consistency with the observed UHECR composition, and discuss a possible enhancement factor that would favor the acceleration of heavier elements in TDEs. In Sec. 6, we discuss the broader consistency of our findings with theoretical and observational expectations, identify areas of tension, and outline directions of future study.
@@ 2.1 TDE source model
The observed neutrino-TDE associations share certain common properties: relatively high blackbody luminosities in the optical-UV range, strong dust echoes in the infrared range – which are delayed by days with respect to the blackbody peaks – and the neutrinos arriving with comparable delays as well. X-rays were also observed in all these cases. The dust echoes are typically interpreted as re-processed radiation from higher energies, where the delays are indicative of the size of the dust torus surrounding the source. Considering that the neutrinos time delays are comparable with , and the suggestion that neutrino-emitting TDEs and TDEs with strong dust echoes might be the same population (van Velzen et al., 2021), it was proposed that the IR photons might be the main target for the neutrino production (e.g., Winter & Lunardini, 2023; Yuan & Winter, 2023) - although a contribution from other wavelengths must be present. This requires that the cosmic ray primaries have high enough rigidity to exceed the photo-pion production threshold
@@
where is the target photon temperature, which means that UHECR primaries are required. The emergent toy model picture is then a production region of size given by the size of the dust torus, within which the infrared photons produce a quasi-isotropic background photon field. The UHECRs will be accelerated somewhere within this region – such as winds interacting with each other or with the circum-TDE material, off-axis or choked jets – or even in some more compact regions (even though the model probably does not capture all contributions to the neutrino production then).
@@
We adopt the quasi-isotropic emission model in Winter & Lunardini (2023) (model M-IR), which uses the measured luminosities and temperatures as closely as possible. The interesting photon targets are the blackbody optical-UV emission (OUV, temperature ), the time-delayed, modelled dust echo in the infrared range (IR, temperature ), and X-rays probably from the innermost regions (temperature ). These photon spectra are modeled by black (gray) body spectra with the individual TDE parameters and time evolutions; see Fig. 3 (black dashed-dotted curves) for two examples. From the (observed) time delays of the IR versus OUV peaks, we obtain the size of the dust torus ; within this radius quasi-isotropic IR photons are expected. Depending on the size of the region , on the magnetic field , and on the cosmic-ray energy , this volume can be calorimetric to CRs – i.e., magnetically confined CRs lose their energy by radiation losses – or even optically thick to them.
@@
We assume that cosmic-ray nuclei ZA are accelerated in an acceleration zone which we do not model explicitly, but instead (similar to AGN blazar models) parameterize by its injection luminosity and maximal rigidity , where . We parameterize the cosmic-ray injection spectrum (density differential in energy, time and volume) as
@@
The non-thermal injection luminosity of nuclei with mass number is assumed to follow the mass accretion rate as
@@
where we use as fiducial parameter;111As a consequence, assumes the role of the “baryonic loading”, a parameter commonly used in GRB or AGN models. Note that this formula assumes the injection of a single nuclear species (); later we will use mixed injection compositions whose total luminosities adds up to . we will later discuss its degeneracy with other parameters in Sec. 6. The mass accretion rate is assumed to have a peak value of and to trace the observed OUV luminosity time evolution. The minimal energy with , which means that all non-thermal nuclei are picked up (injected) at the same minimal gamma factor (or minimal rigidity) – which is relevant if the mass fraction of the progenitor is to be translated into the non-thermal contribution. Note that the maximal ridigity can be directly converted into the maximal injection energy for a single injection isotope. Later, we will, however, combine different injection isotopes under the assumption of equal , which is the simplest possible assumption. This assumption is obtained if the maximal energy is limited by the size of the accelerator (confinement condition from balancing Lorentz and centrifugal forces). If other processes (such as synchrotron losses or photodisintegration) limit the maximal energy, they will change this relationship, and it may not apply anymore.
@@
In summary, our key parameters, which we vary, are: injection isotope ZA, maximal rigidity and radius of the production region . The studied ranges and are chosen to cover the parameters of Winter & Lunardini (2023) for all three TDEs, and to satisfy the constraints from the non-observation of electromagnetic cascades of hadronic origin (Yuan & Winter, 2023). The maximal rigidity, which earlier was an ad hoc choice beyond the photo-pion threshold of IR photons (Winter & Lunardini, 2023), is varied in a reasonable range to potentially describe UHECR data, see e.g. Heinze et al. (2019). The quasi-thermal photon temperatures, their densities, and their time evolutions are derived from observations of the individual TDEs, as described in detail in Winter & Lunardini (2023). We fix the magnetic field to be for the sake of simplicity, as in Winter & Lunardini (2023). A summary of the model’s parameter values is provided in Tab. 1.

### v3

@@ Abstract
We revisit the Ultra-High-Energy Cosmic Ray (UHECRs) production in Tidal Disruption Events (TDEs) in the light of recent neutrino-TDE associations. We use an isotropically emitting source-propagation model, which has been developed to describe the neutrino production in AT2019dsg, AT2019fdr, and AT2019aalc. These TDEs have strong dust echoes in the infrared range, which are potentially linked with the neutrino production. A mechanism where neutrinos originate from cosmic ray scattering on infrared photons implies cosmic rays in the ultra-high energy range, thus suggesting a natural connection with the observed UHECR. We extrapolate the three TDE associations to a population of neutrino- and UHECR-emitting TDEs, and postulate that these TDEs power the UHECRs. We then infer the source composition, population parameters, and local rates that are needed to describe UHECR data. We find that UHECR data point towards a mix of light to mid-heavy injection isotopes, which could be found, e.g., in oxygen-neon-magnesium white dwarfs, and to a contribution of at least two groups of TDEs with different characteristics, dominated by AT2019aalc-type events. The required local TDE rates of , however, are more indicative of the disruption of main sequence stars. We propose an enhanced efficiency in the acceleration of heavier nuclei that could address this discrepancy. The predicted diffuse neutrino fluxes suggest a population of astrophysical neutrino sources that can be observed by future radio neutrino detection experiments. The derived source parameters are consistent with those expected from the individual neutrino observations.
@@ 1 Introduction
Tidal Disruption Events (TDEs) are astrophysical transients originating from stars that are tidally destroyed by a supermassive black hole (SMBH). The accretion of bound stellar materials could power a month to year-long flare (M. J. Rees, 1988; E. S. Phinney, 1989), whose spectrum extends from radio to X-ray frequencies. There have currently been 170 TDEs, TDE candidates and TDE suspects (S. van Velzen et al., 2011; E. Hammerstein et al., 2022; M. Masterson et al., 2024). They have been identified by the Zwicky Transient Facility (ZTF, E. C. Bellm et al., 2019), the Wide-field Infrared Survey Explorer (WISE, E. L. Wright et al., 2010), and X-ray/radio surveys, such as eROSITA (S. Sazonov et al., 2021) and Very Large Array Sky Survey (VLASS, M. Lacy et al., 2020). Remarkably, the correlation study of an 200 TeV IceCube neutrino event IC-191001A and optical transients has revealed that a TDE, AT2019dsg, could be its most likely astrophysical counterpart (R. Stein et al., 2021). The subsequent joint analyses of IceCube neutrino events and electromagnetic signals have discovered more TDE candidates with potential neutrino associations, including AT2019fdr (S. Reusch et al., 2022), AT2019aalc (S. van Velzen et al., 2021), two dust-obscured TDE candidates (N. Jiang et al., 2023), and AT2021lwx (located at a relatively high redshift , C. Yuan et al., 2024a).
@@
The potential correlations between TDEs and astrophysical neutrinos provide the smoking-gun evidence of cosmic ray accelerations since high-energy neutrinos are generated from the interactions between energetic nuclei and ambient radiation/matter backgrounds. However, for TDEs, the sites for particle accelerations and interactions are still unclear. Many models including relativistic jets (X.-Y. Wang et al., 2011; X.-Y. Wang & R.-Y. Liu, 2016; L. Dai & K. Fang, 2017; C. Lunardini & W. Winter, 2017; N. Senno et al., 2017), accretion disks (K. Hayasaki & R. Yamazaki, 2019), wide-angle outflows/hidden winds (K. Fang et al., 2020), and tidal stream interactions (L. Dai et al., 2015; K. Hayasaki & R. Yamazaki, 2019) have been proposed as the origin of the non-thermal electromagnetic (EM) and neutrino emission from TDEs, which could potentially explain these TDE-neutrino coincidences (e.g., R.-Y. Liu et al., 2020; K. Murase et al., 2020; W. Winter & C. Lunardini, 2021; H.-J. Wu et al., 2022; K. Hayasaki, 2021; W. Winter & C. Lunardini, 2023; C. Yuan & W. Winter, 2023).
@@
The non-detection of -rays from these neutrino-coincident TDEs by the Fermi Large Space Telescope (LAT) exerts stringent constraints on the TDE source parameters. In particular, the recent investigation of electromagnetic radiation induced by secondary particles of hadronic (e.g., and ) processes demonstrates that the production of Ultra-High-Energy Cosmic Rays (UHECRs) with energies eV is permitted in an isotropic radiation region of radius cm without exceeding the -ray upper limits (C. Yuan & W. Winter, 2023). Moreover, it was observed that the neutrino-TDE associations are accompanied by strong dust echoes in the infrared wavelength band, and that the arrival times of the coincident neutrinos were close to the peak of the infrared luminosity (which is delayed with respect to the OUV blackbody peak). This led to proposing cosmic ray scattering on infrared photons as a primary neutrino production mechanism, where cosmic-ray primary energies in the UHECR range are needed to exceed the photopion production threshold (e.g., W. Winter & C. Lunardini, 2023; C. Yuan & W. Winter, 2023; C. Yuan et al., 2024a). These conclusions support the idea that TDEs can be promising UHECR sources, and thus offer a possible answer to the long-standing problem of the origin of the observed UHECR flux (G. R. Farrar & A. Gruzinov, 2009; G. R. Farrar & T. Piran, 2014; B. T. Zhang et al., 2017; C. Guépin et al., 2018; D. Biehl et al., 2018b; T. Piran & P. Beniamini, 2023).
@@
The origin of UHECRs is another puzzle in astroparticle physics (R. Alves Batista et al., 2019). It is challenging to associate the measured CRs with a specific source class, (F. Capel & D. J. Mortlock, 2019; A. Abdul Halim et al., 2024a; G. R. Farrar, 2014; V. Shaw et al., 2022). Even with theoretical modeling of promising source classes, it remains unclear which specific source class is responsible for UHECRs, including Gamma-Ray Bursts (GRBs, J. Heinze et al., 2020; D. Boncioli et al., 2019; R.-Y. Liu et al., 2011; K. Murase et al., 2006; F. Samuelsson et al., 2019; M. Vietri, 1995; E. Waxman, 1995) , Active Galactic Nuclei (AGNs, X. Rodrigues et al., 2018; E. Resconi et al., 2017; C. D. Dermer & S. Razzaque, 2010), Starburst galaxies (Romero, G. E. et al., 2018; L. A. Anchordoqui, 2018; L. A. Anchordoqui et al., 1999), and Galaxy clusters (H. Kang et al., 1997; K. Murase et al., 2008; K. Kotera et al., 2009). Sources that produce UHECRs also might emit neutrinos, which contribute to the high-energy diffuse neutrino background. Therefore, the joint consideration of UHECR data and the IceCube diffuse neutrino background would provide valuable insights into the physical environments (e.g., the abundance of intermediate and heavy elements), the acceleration mechanisms (e.g., how the acceleration efficiency depends on the nuclei mass), and redshift distributions of UHECR sources.
@@
In this study, we revisit the potential role of TDEs as sources of UHECRs, motivated by the recent identification of a possible link between three observed TDEs/candidates (AT2019dsg, AT2019fdr, and AT2019aalc) and high-energy neutrino detections by IceCube. While AT2019dsg is well established as a TDE, the classification of AT2019fdr and AT2019aalc remains uncertain, since both occurred in active galaxies where AGN activity can mimick tidal disruption flares; hence, these events are often more broadly referred to as “accretion flares” (S. van Velzen et al., 2021). In addition, the interpretation of AT2019fdr as a superluminous supernova could also account for the coincident neutrino detection (T. Pitik et al., 2022). Based on the models presented in W. Winter & C. Lunardini (2023), we aim to explore the range of parameters that describe TDEs by analyzing various factors, including the type of disrupted stars, the size of the production region, and the maximum energies of CRs. By examining these parameters, we aim to identify scenarios that reproduce the UHECR data and satisfy bounds from the diffuse neutrino flux observations. To achieve this, we develop detailed computational models of cosmic ray emission in TDEs using the NeuCosmA (P. Baerwald et al., 2011; S. Hümmer et al., 2010) software and then model the transport of UHECRs from sources to Earth using the PriNCe code (J. Heinze et al., 2019). By comparing the results with the Akaike Information Criterion, we determine which scenarios best fit current observational data, leading to an improved understanding of TDEs as potential sources of UHECRs.
@@
The paper is structured as follows: In Sec. 2, we develop a model of TDEs using AT2019dsg, AT2019fdr, and AT2019aalc as templates. We identify a set of parameters that will be used in a fit to the data, and determine the ranges of these parameters that are allowed by the observations of these TDEs and TDE candidates. In Sec. 3, we review the current understanding of TDEs as a population of sources, and introduce our scenarios of reference of the UHECR-neutrino-emitting population of TDEs. Using a fitting procedure, we compare these scenarios with the data on the UHECR spectrum and composition from Auger, and present a detailed analysis of the best-fit scenario and how well it compares to the data. In Sec. 5 we elaborate on the question of the consistency with the observed UHECR composition, and discuss a possible enhancement factor that would favor the acceleration of heavier elements in TDEs. In Sec. 6, we discuss the broader consistency of our findings with theoretical and observational expectations, identify areas of tension, and outline directions of future study. Throughout the paper, we use CGS units unless otherwise specified.
@@ 2.1 TDE source model
The observed neutrino-TDE associations share certain common properties: relatively high blackbody luminosities in the optical-UV range, strong dust echoes in the infrared range – which are delayed by days with respect to the blackbody peaks – and the neutrinos arriving with comparable delays as well. X-rays were also observed in all these cases. The dust echoes are typically interpreted as re-processed radiation from higher energies, where the delays are indicative of the size of the dust torus surrounding the source. Considering that the neutrino’s time delays are comparable with , and the suggestion that neutrino-emitting TDEs and TDEs with strong dust echoes might be the same population (S. van Velzen et al., 2021), it was proposed that the IR photons might be the main target for the neutrino production (e.g., W. Winter & C. Lunardini, 2023; C. Yuan & W. Winter, 2023) - although a contribution from other wavelengths must be present. This requires that the cosmic ray primary nuclei of mass number and charge number have high enough energies to exceed the photo-pion production threshold. The (kinematical) photopion threshold is given by , where is the nucleon energy, is the target photon energy, is the proton mass, is the pion mass, is the speed of light; here heads-on collisions are assumed at threshold.
In a superposition ansatz, where the energy per nucleon , the threshold condition can be translated into one for the rigidity , where is the nuclear momentum and is the elementary charge:
Thus, for heavy isotopes, where in the stability valley, one needs for target photons in the infrared range, which is independent of the composition to a first approximation. Consequently, the maximal rigidity (which we will use instead of the maximal energy) has to exceed this value, which means that the primary cosmic rays will be in the UHECR energy range. Since the neutrino energy , their interaction will produce neutrinos in excess of about 35 PeV. The emergent discrepancy between the high neutrino energies from UHECR interactions and the observed (mean) neutrino energies below a PeV is a challenge applicable to many source classes for which an UHECR connection is postulated; an exception are sources with strong (kG) magnetic fields such as Gamma-Ray Bursts, where synchrotron cooling of the secondary muons, pions and kaons can break the direct connection between UHECR and neutrino energies, see e.g. W. Winter (2012); we will return to this issue in Sec. 6.
@@
The emergent toy model picture is then a production region of size given by the size of the dust torus, within which the infrared photons produce a quasi-isotropic background photon field. The UHECRs will be accelerated somewhere within this region – such as winds interacting with each other or with the circum-TDE material, off-axis or choked jets – or even in some more compact regions (even though the model probably does not capture all contributions to the neutrino production then).
@@
We adopt the quasi-isotropic emission model in W. Winter & C. Lunardini (2023) (model M-IR) which uses the measured luminosities and temperatures as closely as possible. The interesting photon targets are the blackbody optical-UV emission (OUV, temperature with being the Boltzmann constant), the time-delayed, modelled dust echo in the infrared range (IR, temperature ), and X-rays probably from the innermost regions (temperature ). These photon spectra are modeled by black (gray) body spectra with the individual TDE parameters and time evolutions, see Fig. 3 (black dashed-dotted curves) for two examples. From the (observed) time delays of the IR versus OUV peaks we obtain the size of the dust torus ; within this radius quasi-isotropic IR photons are expected. Depending on the size of the region (inferred from ), on the magnetic field , and on the cosmic-ray energy , this volume can be calorimetric to CRs – i.e., magnetically confined CRs lose their energy by radiation losses – or even optically thick to them.
@@
We assume that cosmic-ray nuclei ZA are accelerated in an acceleration zone which we do not model explicitly, but instead (similar to AGN blazar models) parameterize it by its injection luminosity and maximal rigidity , where . The cosmic-ray injection spectrum (density differential in energy, time and volume) is assumed as
@@
where is defined as the fraction of the energy density injected as element above and is a normalization factor that will be derived from Eq. (3) below. The non-thermal injection luminosity of nuclei with mass number is assumed to follow the mass accretion rate as
@@
where we use as fiducial parameter;111As a consequence, assumes the role of the “baryonic loading”, a parameter commonly used in GRB or AGN models. Note that this formula assumes the injection of a single nuclear species (); later we will use mixed injection compositions whose total luminosities adds up to . we will later discuss its degeneracy with other parameters in Sec. 6. The mass accretion rate is assumed to have a peak value of 222The mass accretion rate is obtained from order-of-magnitude estimations (e.g., C. Yuan & W. Winter, 2023). Noting that during the TDE peak accretion phase, the accretion rate could be super-Eddington, e.g., (L. Dai et al., 2018), where is the Eddington accretion rate that can be related to the Eddington luminosity via using the radiation efficiency . Therefore, we estimate , and adopt the fiducial value for consistency with previous studies. and to trace the observed OUV luminosity time evolution. The minimal energy with , which means that all non-thermal nuclei are picked up (injected) at the same minimal gamma factor (or minimal rigidity) – which is relevant if the mass fraction of the progenitor is to be translated into the non-thermal contribution. Note that the maximal ridigity can be directly converted into the maximal injection energy for a single injection isotope. Later, we will, however, combine different injection isotopes under the assumption of equal , which is the simplest possible assumption. This assumption is obtained if the maximal energy is limited by the size of the accelerator (confinement condition from balancing Lorentz and centrifugal forces). If other processes (such as synchrotron losses or photodisintegration) limit the maximal energy, they will change this relationship, and it may not apply anymore.
@@
In summary, our key parameters, which we vary, are: injection isotope ZA, maximal rigidity and radius of the production region . The studied ranges and are chosen to cover the parameters of the original study for all three TDEs, and to satisfy the constraints from the non-observation of electromagnetic cascades of hadronic origin (C. Yuan & W. Winter, 2023). The maximal rigidity, which earlier was an ad hoc choice beyond the photo-pion threshold of IR photons (W. Winter & C. Lunardini, 2023), is varied in a reasonable range to potentially describe UHECR data, see e.g. J. Heinze et al. (2019). The quasi-thermal photon temperatures, their densities, and their time evolutions are derived from observations of the individual TDEs, as described in detail in W. Winter & C. Lunardini (2023). We fix the magnetic field to be for the sake of simplicity, as in W. Winter & C. Lunardini (2023). A summary of the model’s parameter values is provided in Tab. 1.